OpenStudy (anonymous):

A set of equations is given below: Equation A: y = x + 1 Equation B: y = 4x + 5 Which of the following steps can be used to find the solution to the set of equations? x + 1 = 4x + 5 x = 4x + 5 x + 1 = 4x x + 5 = 4x + 1

OpenStudy (anonymous):

@iGreen i didn't mean to close it

OpenStudy (jack1):

Equation A: y = x + 1 Equation B: y = 4x + 5 its just logic man if y = x+1 and y = 4x + 5, then x+1 = 4x + 5 (as y = y)

OpenStudy (igreen):

Already figured it out: http://openstudy.com/study#/updates/551bdea9e4b05e8bcb034ae3

OpenStudy (anonymous):

Thank you

OpenStudy (igreen):

No problem.